Xlogmaster FAQ

Frequently asked questions about the Xlogmaster Last Modified: 24.8.1998

Georg C. F. Greve <greve@gnu.org>

1 1. Introductory Questions

1.1 1.1. Preface

This is the introduction to a list of frequently asked questions (FAQ’s) about the Xlogmaster with answers. This article contains a listing of the questions; subsequent articles contain the questions and answers.

If you know the answer of a question is in the FAQ list, please reply to the question by e-mail instead of posting. Help reduce noise!

Please suggest new questions, answers, wording changes, deletions, etc. The most helpful form for suggestions is a context diff (i.e. the output of ‘diff -c’). Include ‘FAQ’ in the subject of messages sent to me about the FAQ list.

Please do not send questions to me just because you do not want to disturb a lot of people and you think I would know the answer. I do not have time to answer questions individually and keep up with everything else I have to get done. Questions to me that have been answered before slow down development and that is bad for everyone.

1.2 1.2. What is the Xlogmaster ?

The Xlogmaster is a GTK+ (see the GIMP tookit page under http://www.gimp.org/ for more information) based program that helps you to monitor all kinds of system activity. It allows monitoring and background surveillance of all logfiles and devices that allow being read by cat like the ‘/proc’ devices. You can configure the Xlogmaster graphically to take almost any action you can think of upon certain events and it can certainly help you keeping track of everything and increasing system security.

1.3 1.3. What is the current version of the Xlogmaster ?

The current version is 1.4.3, released on August 24. 1998.

1.4 1.4. Where can I find it?

The source can be found via anonymous FTP at ‘ftp.gnu.org:/pub/gnu/’.

1.5 1.5. Where can I get help for using the Xlogmaster ?

If you installed everything correctly the easiest ways are probably

• the tutorial that should be in your info tree. Otherwise you can find it in the ‘doc’ directory of the distribution

• the manpage.

• executing the Xlogmaster via xlogmaster --help

Other ways are to check out the homepage at

          ‘http://www.gnu.org/software/xlogmaster/xlogmaster.html’

or the mailinglists

•  ‘beta-xlogmaster@gnu.org’

Subscription is done by sending a mail to

          ‘beta-xlogmaster-request@gnu.org’

leave the subject blank and enter ‘subscribe’ in the body.

•  ‘help-xlogmaster@gnu.org’

Subscription is done by sending a mail to

          ‘help-xlogmaster-request@gnu.org’

leave the subject blank and enter ‘subscribe’ in the body.

2 2. Compiling the Xlogmaster

2.1 2.1. Normal installation

Normal installation is done by unpacking the archive and then entering it. ./configure ; make ; make install should do everything then. The default is to install it into the ‘/usr/local’ structure. If you want to change that and for other configuration options,

please see ./configure --help.

2.2 2.2. Make doesn’t seem to understand the Makefiles

This can happen on some systems (especiall SunOS or old Solaris ones) – the problem is easily fixed by using gmake instead of make.

Table of Contents

1 1. Introductory Questions . . . . . . . . . . . . . . . . . . 1

1.1 1.1. Preface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 1.2. What is the Xlogmaster ?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 1.3. What is the current version of the Xlogmaster ? . . . . . . . 1 1.4 1.4. Where can I find it? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.5 1.5. Where can I get help for using the Xlogmaster ?. . . . . . . 1

2 2. Compiling the Xlogmaster . . . . . . . . . . . . . . . 2

2.1 2.1. Normal installation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 2.2. Make doesn’t seem to understand the Makefiles . . . . . . . 2